# Generating metamers of human scene understanding

### v1

@@ 1 Introduction
Understanding the latent representation of a scene formed by humans after viewing remains a fundamental unanswered challenge in cognitive science (Epstein and Baker, 2019; Bonner and Epstein, 2021; Malcolm et al., 2016; Võ, 2021). What is clear is that humans represent coherent scenes by a mixture of “gist” information encoded from peripheral vision (Potter, 1975; Greene and Oliva, 2009) with high-resolution but sparse information that humans extract during their scene viewing fixations (Larson and Loschky, 2009; Larson et al., 2014; Eberhardt et al., 2016). Related recent work on scene perception has focused on the concept of object and scene metamers—generated stimuli that, although physically different from originals, cannot be discriminated as different by humans when viewed under constrained experimental conditions. (Freeman and Simoncelli, 2011; Balas et al., 2009; Rosenholtz et al., 2012). Understanding scene metamerism is important because metamers tell us the level of misalignment between an actual and generated image that is tolerated by humans and judged to be the same. Generated scenes that fail to become metamers also reveal the details that are important to a scene’s representation and that, if changed, result in the generation being detected. However, although several paradigms have been used to identify metamers (e.g., same-different tasks, A/B/X tasks, oddity judgment tasks) (Rosenholtz, 2020), this work on scene perception used simple generative models to synthesize textures and shapes that were shown in behavioral experiments to be metameric with what humans perceive in their visual periphery when the eye position is fixed. These paradigms, however, were not designed to study how post-gist changes in fixation affect scene metamerism or what objects a person believes to exist in their blurred peripheral view of a scene, which are the problems we engage.
@@
Inspired by these previous studies showing that generated textures and shapes can become metamers for human scene perception, we introduce MetamerGen, a state-of-the-art generative model that extends the metamer generation approach to human scene understanding. Rather than seeking to generate simple patterns that share low-level statistics with peripheral vision, MetamerGen better captures a post-gist level of representation reflecting multiple free-viewing fixations. We see this topic as closer to scene understanding because we are seeking to generate a hypothesis for what a person believes to be in their peripheral vision, and henceforth we will use the term scene metamer to refer to two scenes that have an equivalent understanding. Our approach combines a gist-level scene representation extracted from peripherally blurred pixels with higher-resolution and fixation-specific “foveal” representations corresponding to scene-viewing fixations. Scene gist and the objects fixated during viewing are therefore used to generate in the non-fixated blurred pixels a scene context that is aligned with what a human understands to be in their peripheral vision.
@@
To adapt the Stable Diffusion model to our task of generating a scene in blurred peripheral pixels, we introduce a dual-stream representation of foveated scenes (i.e., ones with a high-resolution center and blurred periphery) using a self-supervised image encoder (DINOv2) (Caron et al., 2021; Oquab et al., 2024; Darcet et al., 2024). We utilize an adapter-based framework (Mou et al., 2023), where we condition a pre-trained text-to-image diffusion model on fixation-grounded features extracted by DINOv2 feature representations obtained at each of the fixation locations. We complement the fixation representations with peripheral information, adding a second source of conditioning that uses DINOv2 tokens extracted from a blurred-out version of the same image, capturing the context.
Our conditioning mechanism allows us to generate plausible scene hypotheses from a variable information input, where more foveal glimpses of a scene will lead to a richer DINOv2 representation that enables MetamerGen to generate increasingly plausible and contextually appropriate content at the non-fixated scene locations, analogous to how human scene understanding becomes more elaborate with more viewing fixations. We see MetamerGen as a tool for generating fixation-specific scene understanding hypotheses that cognitive scientists can test in behavioral studies.
@@ 2.1 Image generation using latent generative models
Latent diffusion models (LDMs) (Rombach et al., 2022) reduce the overall cost by applying the diffusion processes in the latent space of a variational autoencoder (VAE) (Kingma and Welling, 2013). Stable Diffusion (Rombach et al., 2022) uses a pre-trained VAE that spatially compresses images $8\times$ with its encoder and decompresses latent diffusion samples with the corresponding decoder. The denoiser $\epsilon_{\theta}(\cdot)$ is a UNet (Ronneberger et al., 2015) consisting of pairs of down and up-sampling blocks at four resolution levels, as well as a middle bottleneck block. Each network block consists of ResNet (He et al., 2015), spatial self-attention, and cross-attention layers, with the latter introducing the conditioning information.
@@
$Q=FW_{Q},\ K=eW_{K},\ V=eW_{V}$
$Q\in\mathbb{R}^{hw\times d_{k}},\ K\in\mathbb{R}^{n\times d_{k}},\ V\in\mathbb{R}^{n\times d_{v}}$ (1)
@@ 2.2 Self-supervised image encoders
DINOv2 (Caron et al., 2021; Oquab et al., 2024) is a self-supervised vision transformer trained for hierarchical visual representation learning without manual annotations. Using multiple self-supervised objectives, including a contrastive loss that causes image features that appear together to have similar embeddings and a reconstruction loss that induces patches to redundantly encode information about their surrounding context, DINOv2 represents both local visual details and higher-level semantics. These properties make it an excellent tool to study fixation-by-fixation human scene understanding. Adeli et al. (2023; 2025) have shown how self-supervised encoders were capable of capturing object-centric representations without labels as well as providing a backbone capable of predicting high-level neural activity in the brain.
@@ 2.3 Adapting latent diffusion models to new conditions
3 Perceptually-informed conditioning
Figure 1: MetamerGen model architecture. High-resolution and blurred low-resolution images are processed through DINOv2-Base to extract patch tokens each. Foveal features are obtained by applying binary masks to high-resolution patch tokens, retaining only fixated regions. Both foveal and peripheral patch tokens are processed through separate Perceiver-based query networks that compress features into conditioning tokens compatible with Stable Diffusion’s cross-attention mechanism. The resulting dual conditioning streams are integrated into the pretrained UNet for guided image denoising and generation.
@@ 3.1 Representing foveal & peripheral visual features
Given an image and a set of fixation locations, potentially made by a human during free-viewing, we first aim to extract the foveal information from the fixation locations and the peripheral information regarding the overall image context. We employ a DINOv2-Base model (with registers) as the feature extractor to obtain these two sources of information. In Appendix Section A.10 we validate the choice of DINOv2 as the feature extractor for MetamerGen by showing its superiority to CLIP.
DINOv2 processes $448\times 448$ images with a patch size of $14\times 14$, yielding $1024$ tokens ($32\times 32$ grid), each embedded in $768$ dimensions (along with a CLS token representing the entire image, and four register tokens encoding general information about the image). The patch token at a specific location encodes detailed visual and semantic information about that location, analogous to the high-resolution information sampled by the fovea during a fixation. It also encodes limited information about the location’s visual context, analogous to low-resolution parafoveal information (Adeli et al., 2023; 2025). To model the information gathered during a series of fixations, we apply a binary mask $M_{\text{fixation}}$ to the patch tokens extracted from a scene image $I$, corresponding to the image locations fixated by humans, zeroing out all non-fixated image patches.
For peripheral visual features, we simulate the inherent uncertainty in peripheral vision by downsampling the input image, and then upsampling it back to $448\times 448$. The blurry image, $I_{\text{peripheral}}$, is also processed with DINOv2, but now retaining all output patch tokens without masking. These peripheral tokens encode uncertain visual representations across the entire scene, capturing the noisy information available in peripheral vision that requires validation through targeted foveal fixations (Srikantharajah and Ellard, 2022; Michel and Geisler, 2011).
@@ 3.2 Foveal & peripheral conditioning adapters
Both foveal and peripheral DINOv2 embeddings are first processed through separate Perceiver-based resampler networks $R(\cdot)$ (Alayrac et al., 2022; Jaegle et al., 2021) that compress the 1024 DINOv2 embeddings into 32 conditioning tokens compatible with the pre-trained UNet’s cross-attention. (For more information, please refer to Appendix A.3).
@@
$K_{c}=e_{c}W_{K}^{c},\ V_{c}=e_{c}W_{V}^{c},$
$K_{c}\in\mathbb{R}^{n_{c}\times d_{k}},\ V_{c}\in\mathbb{R}^{n_{c}\times d_{k}},\ c=\{\text{text},\ \text{foveal},\ \text{peripheral}\}$ (4)
@@
$\text{Attention}(Q,K,V)=\$$\text{softmax}\left(\frac{QK_{\text{text}}^{T}}{\sqrt{d_{k}}}\right)V_{\text{text}}+\lambda_{\text{foveal}}\cdot\text{softmax}\left(\frac{QK_{\text{foveal}}^{T}}{\sqrt{d_{k}}}\right)V_{\text{foveal}}$
$+\lambda_{\text{peripheral}}\cdot\text{softmax}\left(\frac{QK_{\text{peripheral}}^{T}}{\sqrt{d_{k}}}\right)V_{\text{peripheral}}$ (5)
$\lambda_{\text{foveal}}$ and $\lambda_{\text{peripheral}}$ are scaling factors that control the contribution of of either foveal or peripheral visual features to the generation process. In practice we “freeze” the text conditioning, by setting the text caption for all images to an empty string “ ”.
@@ 3.3 Training and Inference
To enable robust conditioning during inference, we randomly drop conditions with probabilities $p_{\text{foveal}}=0.05$ and $p_{\text{peripheral}}=0.10$. The higher peripheral dropout rate prevents over-reliance on peripheral features, which despite blurred image features retain substantial visual information compared to the sparse foveal features. We employ the DDIM sampler (Song et al., 2022) for $50$ timesteps, with CFG++ (Chung et al., 2025). We set $\lambda_{\text{foveal}}=1.2$ and $\lambda_{\text{peripheral}}=0.7$ to balance detail generation with scene plausibility.
We point out that although MetamerGen is conditioned on dense DINOv2 representations of an image (periphery and fixation DINOv2 patch embeddings), the model does not simply reconstruct input images verbatim. We attribute this to the lossiness introduced by the DINOv2 embeddings, as well as stochasticity in the sampling process. We demonstrate this further in Appendix A.7 and A.8.
4 MetamerGen for image generation
@@
We first evaluate the image quality of samples from our model using Fréchet Inception Distance (FID; Heusel et al., 2017) between images generated from MetamerGen and COCO-10k-test. Figure 2 shows the results using a single central fixation. Green: we fix the blur level to $0.25\times$, matching our behavioral paradigm, and evaluate how peripheral context affects generation quality by varying the peripheral scale. As peripheral scale increases, FID scores improve showing that the model is able to better integrate the context coming from the peripheral DINOv2 representations. Red: we evaluated the effect of the blur level, showing that our model can consistently generate plausible scenes for all levels of blur. We include a text-to-image baseline ( Blue) using SD-1.5 with 10k random captions from the COCO training set. MetamerGen, fine-tuned on the COCO images, consistently outperforms the text-to-image model, proving that we have successfully integrated images of variable resolution into the conditioning mechanism of Stable Diffusion.
@@ 5 Behaviorally-conditioned scene metamers
5.1 Probing latent scene representations through metamer judgments
Initially established in color science to reveal trichromatic vision, metamers have since been applied to texture perception and visual crowding to infer the underlying structure of human perceptual systems. In the context of scene understanding, metamers offer a unique opportunity to probe what the visual system extracts and retains from complex natural scenes.
@@ 5.2 Real-time behavioral paradigm
Figure 3: Metameric vs. non-metameric judgments. (Left) Original images with human fixations overlaid in red and corresponding generated images judged as ”same” by participants. (Right) Original images with fixations and generated images judged as ”different” by participants. More examples based off of both human-fixation and random-fixation guided generations can be seen in Appendix A.6
@@
We employed 45 participants in a naturalistic free-viewing same-different paradigm. Each trial followed a structured sequence (Appendix  A.1.2): participants first completed a drift check, fixated on a central cross, then freely viewed a natural scene image until reaching a predetermined fixation count $\{1,2,3,5,10\}$, after which the image automatically disappeared. Critically, participants chose their own fixation locations (see Appendix A.1.1). We systematically varied information availability by manipulating fixation count, testing how additional visual information influenced the generation quality.
During a subsequent 5-second interval, participants maintained central fixation while our system processed their actual fixation coordinates and the original image in real-time. MetamerGen’s generated image then appeared briefly for 200 milliseconds—too brief to allow eye movements but sufficient for perceptual comparison (Broderick et al., 2023; Wallis et al., 2019). Participants used a gamepad to indicate whether this second image matched their initial percept.
Participants encountered two primary experimental conditions: metamers generated from their own fixations, and identical original images reshown as controls. As a third comparison condition, 12 participants additionally saw metamers generated from randomly-sampled coordinates instead of their actual fixations. While random metamers often fooled participants, they varied so much that they could not be used to explain human behavior. Generating metamers based on human-fixated locations resulted in better-controlled variability, centered around our best estimate of the participant’s scene understanding.
@@
Our stimulus set comprised 300 images from the Visual Genome dataset (Krishna et al., 2017), specifically sourced from the YFCC100M subset (Thomee et al., 2016) to avoid overlap with COCO training data used in MetamerGen training. We employed DreamSim (Fu et al., 2023) to cluster images in semantic representational space and selected one representative image per cluster to maximize visual diversity. Images were filtered to exclude challenging elements for current diffusion models: human hands, faces, and bodies, as well as clocks, text, and numbers.
6 Multiple levels of visual features drive metamer judgments
MetamerGen is conditioned on actual human fixation sequences, providing a richer and more dynamic model of scene understanding. Because it can generate plausible hypotheses for naturalistic images from both peripheral and foveal information (Figure 3), it also enables analysis of which visual features—ranging from low to high levels—shape metameric judgments. In our first analysis, we compared visual similarity from neurally grounded CNN features with human same–different responses. Although images generated from human fixation and random fixation sequences fooled participants about equally often (29.4% and 27.7% of the time, respectively, $p=0.24$), we found a stark contrast in interpretability between these conditions. For human-fixation-based metamers, higher similarity to the original predicted more “same” judgments. For random-fixation-based metamers, however, high similarity often increased “different” judgments, suggesting that realistic details in non-fixated regions may expose inconsistencies with the viewer’s internal scene representation. We confirmed the same pattern across explicitly defined, interpretable feature hierarchies: features at all levels contributed to explaining human metameric judgments, with the fixation-based effect becoming especially pronounced for high-level semantic features (e.g., DreamSim, CLIP).
Figure 4: Multi-level feature analysis pipeline using neurally-grounded model: (Top) Early, mid, and late network layers serve as proxies for different stages of visual processing from V1 to IT. (Bottom) Results show that as feature similarity increased at different processing stages, the proportion of participants judging generated images as metameric also increased. These effects were clearer when metamers were generated based on fixated locations (salmon) than on randomly-sampled locations (turquoise).

### v2

@@ 1 Introduction
Understanding the latent representation of a scene formed by humans after viewing remains a fundamental unanswered challenge in cognitive science (Epstein and Baker, 2019; Bonner and Epstein, 2021; Malcolm et al., 2016; Võ, 2021). What is clear is that humans represent coherent scenes by a mixture of “gist” information encoded from peripheral vision (Potter, 1975; Greene and Oliva, 2009) with high-resolution but sparse information that humans extract during their scene viewing fixations (Larson and Loschky, 2009; Larson et al., 2014; Eberhardt et al., 2016). Related recent work on scene perception has focused on the concept of object and scene metamers—generated stimuli that, although physically different from originals, cannot be discriminated as different by humans when viewed under constrained experimental conditions. (Freeman and Simoncelli, 2011; Balas et al., 2009; Rosenholtz et al., 2012). Understanding scene metamerism is important because metamers tell us the level of misalignment between an actual and generated image that is tolerated by humans and judged to be the same. Generated scenes that fail to become metamers also reveal the details that are important to a scene’s representation and that, if changed, result in the generation being detected. However, although several paradigms have been used to identify metamers (e.g., same-different tasks, A/B/X tasks, oddity judgment tasks) (Rosenholtz, 2020), this work on scene perception used simple generative models to synthesize textures and shapes that were shown in behavioral experiments to be metameric with what humans perceive in their visual periphery when eye position is fixed. These paradigms were not designed to study how post-gist changes in fixation affect scene metamerism or what objects a person believes to exist in their blurred peripheral view of a scene, which are the problems we engage.
@@
Inspired by these previous studies showing that generated textures and shapes can become metamers for human scene perception, we introduce MetamerGen, a state-of-the-art generative model that extends the metamer generation approach to human scene understanding. Rather than seeking to generate simple patterns that share low-level statistics with peripheral vision, MetamerGen better captures a post-gist level of representation reflecting multiple free-viewing fixations. We see this topic as closer to scene understanding, and we see the scenes generated by MetamerGen being hypotheses for what people believe to be in their peripheral vision. Henceforth, we will therefore use the term scene metamer to refer to two scenes that have an equivalent understanding. Our approach combines a gist-level scene representation extracted from peripherally blurred pixels with higher-resolution and fixation-specific “foveal” representations corresponding to scene-viewing fixations. Scene gist and the objects fixated during viewing are used to generate in the non-fixated blurred pixels a scene context that is aligned with what a human understands to be in their peripheral vision.
@@
To adapt Stable Diffusion to our task of generating a scene in blurred peripheral pixels, we introduce a dual-stream representation of foveated scenes (i.e., ones with a high-resolution center and blurred periphery) using a self-supervised image encoder (DINOv2) (Caron et al., 2021; Oquab et al., 2024; Darcet et al., 2024). We utilize an adapter-based framework (Mou et al., 2023), where we condition a pre-trained text-to-image diffusion model on fixation-grounded features extracted by DINOv2 feature representations obtained at each of the fixation locations. We complement the fixation representations with the scene context available from peripheral information by adding a second source of conditioning that uses DINOv2 tokens extracted from a blurred version of the same image.
Our conditioning mechanism allows plausible scene hypotheses to be generated from variable input information. We expect more foveal glimpses of a scene will lead to a richer DINOv2 representation that will in turn enable MetamerGen to generate increasingly plausible and contextually appropriate content at the non-fixated scene locations, analogous to how human scene understanding becomes more elaborated with more viewing fixations. We see MetamerGen as a tool for generating fixation-specific scene understanding hypotheses that cognitive scientists can test in behavioral studies.
@@ 2.1 Image generation using latent generative models
Latent diffusion models (LDMs) (Rombach et al., 2022) reduce the overall cost by applying the diffusion processes in the latent space of a variational autoencoder (VAE) (Kingma and Welling, 2013). Stable Diffusion (Rombach et al., 2022) uses a pre-trained VAE that spatially compresses images $8\times$ with its encoder and decompresses latent diffusion samples with the decoder. The denoiser $\epsilon_{\theta}(\cdot)$ is a UNet (Ronneberger et al., 2015) with pairs of down and up-sampling blocks, as well as a middle bottleneck block. Each network block consists of ResNet (He et al., 2015), spatial self-attention, and cross-attention layers, with the latter introducing the conditioning information.
@@
$Q=FW_{Q},\ K=eW_{K},\ V=eW_{V},\quad Q\in\mathbb{R}^{hw\times d_{k}},\ K\in\mathbb{R}^{n\times d_{k}},\ V\in\mathbb{R}^{n\times d_{v}}$ (1)
@@ 2.2 Self-supervised image encoders
DINOv2 (Caron et al., 2021; Oquab et al., 2024) is a self-supervised vision transformer trained for hierarchical visual representation learning without manual annotations. Using multiple self-supervised objectives, including a contrastive loss that causes image features that appear together to have similar embeddings and a reconstruction loss that induces patches to redundantly encode information about their surrounding context, DINOv2 represents both local visual details and higher-level semantics. These properties make it an excellent tool to study fixation-by-fixation human scene understanding. Adeli et al. (2023; 2025) showed that self-supervised encoders are capable of capturing object-centric representations without labels and can provide a backbone capable of predicting high-level neural activity in the brain.
@@ 2.3 Adapting latent diffusion models to new conditions
3 MetamerGen: perceptually-informed conditioning
Figure 1: MetamerGen model architecture. High-resolution and blurred, low-resolution images are processed through DINOv2-Base to extract patch tokens. Foveal features are obtained by applying binary masks to high-resolution patch tokens, retaining only fixated regions. Both foveal and peripheral patch tokens are processed through separate Perceiver-based query networks that compress features into conditioning tokens compatible with Stable Diffusion’s cross-attention mechanism. The resulting dual conditioning streams are integrated into the pretrained UNet.
@@ 3.1 Representing foveal & peripheral visual features
Given an image and a set of fixation locations, such as those fixated by a human during free-viewing, we first aim to extract the foveal information from the image at each fixation location and peripheral information capturing the overall context. We employ a DINOv2-Base model (with registers) as the feature extractor to obtain these two sources of information. In Appendix Section A.10 we validate our choice of DINOv2 as the feature extractor for MetamerGen by showing its superiority to CLIP.
DINOv2 processes $448\times 448$ images with a patch size of $14\times 14$, yielding $1024$ tokens ($32\times 32$ grid), each embedded in $768$ dimensions (along with a CLS token representing and four register tokens encoding general information about the image). The patch token at a specific location encodes detailed visual and semantic information about that location, analogous to the high-resolution information sampled by the fovea during a fixation. It also encodes local contextual information from around the fixated locations, analogous to the low-resolution parafoveal information extracted by humans(Adeli et al., 2023; 2025). To model the information gathered during a series of fixations, we apply a binary mask $M_{\text{fixation}}$ to the patch tokens extracted from a scene image $I$, corresponding to the image locations fixated by humans, zeroing out all non-fixated image patches.
To obtain peripheral visual features, we downsample the image and then upsample it back to $448\times 448$. This blurred image, $I_{\text{peripheral}}$, is also processed with DINOv2, but now retaining all output patch tokens without masking. These peripheral tokens encode uncertain visual representations across the entire scene, capturing the noisy information available in peripheral vision that requires validation through targeted foveal fixations (Srikantharajah and Ellard, 2022; Michel and Geisler, 2011).
@@ 3.2 Foveal & peripheral conditioning adapters
Both foveal and peripheral DINOv2 embeddings are first processed through separate Perceiver-based resampler networks $R(\cdot)$ (Alayrac et al., 2022; Jaegle et al., 2021) that compress the 1024 DINOv2 embeddings into 32 conditioning tokens compatible with the pre-trained UNet’s cross-attention (see Appendix A.3 for more information).
@@
$K_{c}=e_{c}W_{K}^{c},\ V_{c}=e_{c}W_{V}^{c},\quad K_{c},\ V_{c}\in\mathbb{R}^{n_{c}\times d_{k}},\ c=\{\text{text},\ \text{foveal},\ \text{peripheral}\}$ (4)
@@
$\begin{split}\text{Attention}(Q,K,V)=\ &\text{softmax}\left(\frac{QK_{\text{text}}^{T}}{\sqrt{d_{k}}}\right)V_{\text{text}}+\lambda_{\text{foveal}}\cdot\text{softmax}\left(\frac{QK_{\text{foveal}}^{T}}{\sqrt{d_{k}}}\right)V_{\text{foveal}}\\ &+\lambda_{\text{peripheral}}\cdot\text{softmax}\left(\frac{QK_{\text{peripheral}}^{T}}{\sqrt{d_{k}}}\right)V_{\text{peripheral}}\end{split}$ (5)
$\lambda_{\text{foveal}}$ and $\lambda_{\text{peripheral}}$ are scaling factors that control the contribution of either foveal or peripheral visual features to the generation process. In practice we “freeze” the text conditioning, by setting the text caption for all images to an empty string “ ”.
@@ 3.3 Training and Inference
To enable robust conditioning in inference, we randomly drop conditions with probabilities $p_{\text{foveal}}=0.05$ and $p_{\text{peripheral}}=0.10$. The higher peripheral dropout rate prevents over-reliance on peripheral features, which, despite being blurred, retain substantial visual information compared to the sparse foveal features. We employ DDIM (Song et al., 2022) for $50$ timesteps, with CFG++ (Chung et al., 2025) and set $\lambda_{\text{foveal}}=1.2$ and $\lambda_{\text{peripheral}}=0.7$ to balance detail generation with scene plausibility.
@@
We point out that, although MetamerGen is conditioned on dense DINOv2 representations of an image (periphery and fixation patch embeddings), the model does not simply reconstruct input images verbatim. We attribute this to the lossiness introduced by the DINOv2 embeddings, as well as stochasticity in the sampling process. We demonstrate this further in Appendices A.7 and A.8.
3.4 Image generation quality
We first evaluate the image quality of samples from our model using Fréchet Inception Distance (FID; Heusel et al., 2017) between images generated from MetamerGen and COCO-10k-test. Figure 2 shows the results using a single central fixation. Green: We fix the blur level to $0.25\times$, matching our behavioral paradigm, and evaluate how peripheral context affects generation quality by varying the peripheral scale. As peripheral scale increases, FID scores improve as the model better integrates the context coming from the peripheral DINOv2 representations. Red: We evaluated the effect of blur level and found that our model consistently generates plausible scenes for all levels of blur in our evaluation. Blue: We include a text-to-image baseline using SD-1.5 with 10k random captions from the COCO training set. MetamerGen, fine-tuned on the COCO images, consistently outperforms the text-to-image model, proving that we have successfully integrated images of variable resolution into the conditioning mechanism of Stable Diffusion.
@@ 4 Behaviorally-conditioned scene metamers
4.1 Probing latent scene representations through same/different judgments
Metamerism, a term initially established in the color vision literature, has since been applied to questions in texture perception and visual crowding to infer underlying perceptual representation (Freeman and Simoncelli, 2011; Balas et al., 2009; Rosenholtz et al., 2012). In the context of scene understanding, metamers offer a unique opportunity to probe what the visual system encodes and retains after viewing complex natural scenes.
@@ 4.2 Real-time behavioral paradigm
Participants (n=45) engaged in a two-stage paradigm that consisted of freely viewing a scene followed by making a same-different judgment about that scene. Each trial followed the same structured sequence (Appendix  A.1.2). In Stage 1, participants completed a drift check and then freely viewed a natural scene image until reaching a predetermined fixation count $\{1,2,3,5,10\}$, after which the image was removed. We systematically varied information availability by manipulating fixation count, testing how additional visual information influenced the generation quality and the rate of metamerism. Critically, participants chose their own fixation locations (see Appendix A.1.1). Stage 2 immediately followed scene viewing offset and began with participants maintaining central fixation throughout a 5-second interval. During this interval, MetamerGen generated in real time a new version of the original image based on the viewing fixations. Participants were then shown a second image for 200 milliseconds—too brief for an eye movement but sufficient for a perceptual judgment (Broderick et al., 2023; Wallis et al., 2019), and asked to indicate whether this second image was the same or different from the image viewed originally. Although our primary experimental manipulation was whether the second image was the original or a generation from the participant’s own fixations, we also had a “random” condition consisting of 12 other participants making a similar same or different judgment, only now between a just-viewed original image and a generation based on a random sampling of fixations in the image. We found that, although random generations did sometimes elicit “same” judgments, generations based on a viewer’s own fixations produced the best estimate of their scene understanding.
Figure 3: Metameric vs. non-metameric judgments. (Left) Original images with human fixations overlaid in red paired with generated images judged as the “same” by participants. (Right) Original images with fixations and generated images judged as “different” by participants. More examples, including generations from random fixations, can be found in Appendix A.6
@@
Our stimuli were 300 images from the YFCC100M dataset (Thomee et al., 2016), a subset of the Visual Genome dataset (Krishna et al., 2017). We chose this dataset to avoid overlap with the COCO dataset used in MetamerGen training. To maximize visual diversity in our 300 image sample, we used DreamSim (Fu et al., 2023) to cluster YFCC100M images in semantic representational space and selected one representative image per cluster. Images were filtered to exclude challenging elements for current diffusion models (e.g., humans, printed text, clocks).
5 Results: Probing the features of scene metamers
As shown in Figure 3, MetamerGen’s conditioning on a viewer’s own fixations enables it to generate rich and plausible hypotheses for a person’s scene understanding. However, from such qualitative results it is difficult to know the features responsible for a generated scene becoming a metamer. We conducted three quantitative analyses to answer this question. First, we compared visual similarity from neurally grounded CNN features with human same–different responses. Our second analysis focused on interpretability by asking which visual features—ranging from low to high levels—shape metameric judgments. Third, we determined the contribution of foveal versus peripheral information by ablating each from the combined model and observing the effect on metamerism rates.